### v2

# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services

Abstract
The demand for high-fidelity test data is paramount in industrial settings where access to production data is largely restricted. Traditional data generation methods often fall short, struggling with low-fidelity and the ability to model complex data structures and semantic relationships that are critical for testing complex SQL code generation services like Natural Language to SQL (NL2SQL). In this paper, we address the critical need for generating syntactically correct and semantically “meaningful” mock data for complex schema that includes columns with nested structures that we frequently encounter in Google SQL code generation workloads. We highlight the limitations of existing approaches used in production, particularly their inability to handle large and complex schema, as well as the lack of semantically coherent test data that lead to limited test coverage. We demonstrate that by leveraging Large Language Models (LLMs) and incorporating strategic pre- and post-processing steps, we can generate realistic high-fidelity test data that adheres to complex structural constraints and maintains semantic integrity to the test targets (SQL queries/functions). This approach supports comprehensive testing of complex SQL queries involving joins, aggregations, and even deeply nested subqueries, ensuring robust evaluation of SQL code generation services, like NL2SQL and SQL Code Assistant services. Our results demonstrate the practical utility of an out-of-the-box LLM (gemini) based test data generation for industrial SQL code generation services where generating realistic test data is essential due to the frequent unavailability of production datasets.
I Introduction
In the fast-paced landscape of modern data systems, particularly within SQL development and testing, the demand for high-quality mock data that accurately mirrors real-world scenarios is more critical than ever [1, 2, 3, 4, 5, 6]. This need is amplified in industrial settings where data is often confidential and/or sensitive. In such environments, direct access to production data is severely restricted, and traditional approaches like production data sampling become time-consuming, infeasible, or violate data privacy requirements. Historically, mock data generation has relied on manual creation, production data sampling or mathematical models [5, 2, 7, 8].
One key challenge for effective test data generation lies in achieving realism and diversity within industrial contexts. Mock data must be sufficiently realistic and diverse to effectively simulate the wide range of scenarios encountered in real-world production systems, including both common and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [9, 10]. Another challenge is supporting test cases with complex data structure, consistency, and data integrity requirements [5, 4, 11, 3]. For industrial applications heavily reliant on structured data types such as JSON, XML, protocol buffers (a.k.a. proto [12]), SQL and Open API schemas, it is paramount that the generated mock data strictly adheres to the specified structure and maintains data integrity and consistency [13, 14]. Specifically, our SQL code generation services utlizie GoogleSQL [15] that supports complex nested data types, like array, struct and proto – making it a representative case for the challenges faced by many modern SQL dialects that support nested data structures. Manual processes are highly susceptible to errors that can compromise the validity of testing and development. Furthermore, scalability and efficiency of generation are critical for enterprise needs, especially when a large amount of test data is needed [16, 17, 10]. As enterprise software systems grow in scale and complexity, the need for large volumes of mock data increases exponentially. Manual and heuristic data generation approaches become impractical and unsustainable in such environments, hindering development velocity and quality assurance efforts.
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as schemas and protocol buffers. The use of LLMs for synthetic data generation is becoming popular in various domains [18, 16, 19, 20, 21, 22, 23, 24], the generation of structured data introduces unique challenges that are particularly relevant to industrial applications. LLMs, when appropriately trained, can provide realistic and privacy-preserving data synthesis. They can generate mock data that more closely approximates real-world textual data distributions and patterns, capturing subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [25, 26]. This is crucial for realistic testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [18, 27, 26, 28]. Given a schema or proto definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. However, tabular data generation via language models remains an active and evolving field, facing challenges such as accurately capturing complex inter-column (pairwise) relationships, mitigating the risk of “hallucinations” or data inconsistencies, and effectively handling rare values and outliers (also see Table II for vanilla LLM evaluation results) [26, 28].
Deep generative models [29, 30, 31] have shown promise in tabular data generation, particularly for categorical, numerical, and text attributes. However, they are fundamentally limited by their focus on flat schemas, neglecting nested structures, and their reliance on training data, rendering them infeasible in scenarios where production data is restricted. In contrast, our LLM-based approach, leveraging gemini [32] out-of-the-box, directly addresses these limitations by generating structurally compliant data for arbitrarily complex, nested schemas without requiring fine-tuning.
However, current LLM-based data generation approaches often fall short in several critical areas. A significant limitation lies in their ability to effectively handle more complex data structures commonly found in our workloads. Specifically, LLMs struggle to accurately generate data that conforms to large (many columns or fields) and complex nested structures. While LLMs can generate data that adheres to basic schemas, accurately generating data that conforms to deeply nested structures or schemas with complex dependencies remains a challenge. This limitation is particularly problematic in testing real-world SQL generation services that rely on intricate data models, where the nuances of nested relationships over very wide tables are critical for realistic testing and validation.
Furthermore, ensuring the semantic validity of the generated data is another crucial challenge that must be addressed to generate truly realistic test data. This is even more critical for testing generated SQL queries (e.g., in NL2SQL systems) on user data. While LLMs might generate data that is syntactically correct and structurally valid, it often lacks the necessary semantic coherence across columns and tables. Most test data generation approaches focus on approximating per-column or pairwise column distributions from the production data [29, 30, 21, 9, 1, 3, 4, 5] leading to illogical combinations of column values, semantically incoherent join tables – violating real-world business rules. Overcoming these limitations, including both structural and semantic challenges, is crucial for fully realizing the potential of LLMs in generating comprehensive and accurate mock data for complex workloads.
In this paper, we address the critical need for generating syntactically correct and semantically “meaningful” mock data for structurally well-defined data formats, focusing on the specific requirements of industrial use cases. We define “meaningful” data as data that effectively supports a wide range of complex test cases for Google SQL code generation services, like NL2SQL and SQL code assistance systems. Robust evaluation of such generation systems requires realistic and complex test data, which are often inaccessible due to privacy or lack of complex data generation capabilities. Services, like NL2SQL, support arbitrarily complex use cases, generating intricate SQL queries involving joins, aggregations, and nested subqueries—scenarios prevalent in real-world applications. The meaningful test data generation should also include the ability to generate data that respects various constraints, such as generating data within a specific date range when the SQL query under test includes a corresponding filter (e.g., WHERE year = 2022). Moreover, we emphasize the importance of maintaining column correlations and data dependencies within a schema, which are often crucial for simulating realistic data behavior. While the ability to generate complex, multi-faceted datasets is essential for thoroughly evaluating these queries, semantic integrity of test data is also critical, in that incorrect or illogical data can lead to misleading test results, failing to expose critical flaws in the generated SQL queries. Without such robust testing, the reliability and trustworthiness of SQL code generation services are severely compromised.
Our key contributions are:
  • LLM-based High-Fidelity and Efficient Test Data Generation: We present a novel approach leveraging LLMs, enhanced with pre- and post-processing, to generate high-fidelity test data compliant with complex and large schemas, including nested column types, addressing the limitations of existing data generation methods in production. Our generation method is significantly more efficient than manual generation used in production where production data access is restricted.
  • Enhanced Test Coverage for Complex SQL Queries: We demonstrate that our generated test data significantly improves test coverage of Google SQL code generation services, where many test cases are disabled due to limited access to production data and the difficulties of manual data generation for large, complex schemas. Our fundamental approach is dialect-agnostic and should be readily applicable to other SQL code generation services that face similar challenges.
  • Semantic Coherence With Test SQL Queries: We emphasize and achieve the generation of semantically meaningful and validated test data. We provide a comparative analysis against production data sampling, showcasing the ability of our generated test data to provide meaningful tests that cover edge cases and semantically relevant predicate columns.
The remainder of this paper is structured as follows. Section 2 discusses real-world SQL code generation services that serve as use cases for our work. Section 3 presents our proposed method for generating high-quality mock data. Section 4 describes our experimental setup and presents the evaluation results. Section 5 discusses related work in the field of data generation. Section 6 provides a discussion of our findings and their implications. Finally, Section 7 concludes the paper.
II SQL code generation services
In this section, we examine the challenges of generating meaningful test data for two real-world GoogleSQL [15] SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
TABLE I: Traditional test data generation methods for real-world SQL code generation services. Test coverage is largely limited by the test data quality, especially when the underlying data structure/schema is large and complex with deep nesting.
SQL Code Assistant NL2SQL
Schema complexity High High
Test (SQL query) complexity Low High
Data generation method Manual Sampling111restricted to a limited set of domains/tables in production
Test coverage Low Low
Table I summarizes test & schema complexities, as well as the traditional test data generation methods used for the services. The underlying data schema for both services are complex with many columns/fields with nested fields (nested proto and array types). The test complexity for NL2SQL is even more complex as the service generates SQL queries spanning over multiple tables across many customer domains. Test cases are submitted by the internal developers and also by crowdsourcing. The test SQL queries for SQL Code Assistant are mostly simple (single-table) selection queries; however, the manually curated test data overlooks complex nested structures from the schema limiting the test cases to SQL queries targeting mostly flat schemas with non-nested fields. The SQL query complexity vary, but most of them use nested fields and subqueries with joins. SQL code generation services require high-fidelity test data that adheres to the complex schema with nested structures that support complex and diverse sets of SQL queries/functions. Additionally, NL2SQL service evaluation requires testing if predicted SQL generates a result set that is semantically coherent with the natural language question. This also means that the test data instances should be realistic and aligned with the context of the posed data question. Unfortunately, production sampling, though anonymized, is restricted to a few accessible domains and tables, resulting in incomplete test scenarios and low test coverage.
TABLE II: Evaluation of test data generation for a large Google Ads production customer table. The table contains 72 columns, 65 of which are nested fields (proto type). Due to data privacy restrictions, production data access and sampling were prohibited. Manual and vanilla LLM generation methods struggled with this moderately complex schema. Our proposed method achieved near-perfect results across all evaluation criteria, demonstrating its ability to generate high-quality test data for wide schemas.
Test Data Generation Methods
Evaluation Criteria Manual Vanilla LLM Our Method
Primary key generated Yes No Yes
Number of fields with correct field names 11/72 62/72 72/72
Total number of table columns generated 11/72 71/72 72/72
Number of nested fields with incorrect level of nesting generated 0/10 36/64 0/64
Number of nested fields with at least one correct field name 10/10 13/64 64/64
Number of nested fields with correct enum values generated 10/10 2/50 50/50
Number of nested fields with correct scalar values generated 10/10 5/64 64/64
Number of nested fields with all the column values generated 9/10 0/64 64/64
Generation time per instance (50p) 1 hour [^2] An engineer, familiar with the table schema, was tasked with generating a test data instance within one hour. While individual speed may vary, it is practically impossible to curate a comprehensive instance in such a limited timeframe. 37 sec 167 sec
II-A SQL Code Assistant
SQL Code Assistant is designed to aid developers in writing SQL queries or SQL functions that access data models (schema, proto). Its testing involves the generated SQL queries/functions and test data instances that are often manually prepared and verified. This manual data generation often leads to incomplete data instances with partially filled schema elements neglecting complex and nested structures. This can lead to low test coverage, limiting the range of executable test queries. Furthermore, even a simple test case over a primitive (char) column can have low coverage if the data instances do not cover all the edge/branch cases under test.
1CREATE PUBLIC FUNCTION GetBalanceInUsd(Input fake_table) 2RETURNS INT64 3AS ( 4 CASE 5 WHEN private_info.running_balance.currency = ’USD’ THEN private_info.running_balance.amount 6 WHEN private_info.running_balance.currency = ’GBP’ THEN private_info.running_balance.amount * 1.26 7 WHEN private_info.running_balance.currency = ’EUR’ THEN private_info.running_balance.amount * 1.05 8 ELSE NULL 9 END 10);
Figure 1: Simple selection query accessing deeply nested fields on a fake table, currency and amount.
Figure 1 illustrates a simple selection query function accessing nested columns (nested proto type fields, fake_table $\rightarrow$ private_info $\rightarrow$ running_valance). The table schema is very large with 161 columns and the column (proto type) that this query is accessing has 3 levels of nesting. The query function is tested using manually curated test data instances that are often partially filled with a few fixed test values on the required fields and/or default values on the nested structures and fields. We found that this leads to low test coverage on many user defined SQL functions for the SQL Code Assistant service. For instance, the above example requires at least three test data instances with all three distinct currency column values, “USD”, “GBP” and “EUR” for the optimal test coverage. As the test complexity increases with more predicates and nesting, providing all test data instances with all the relevant test values becomes much harder. As a result, this was limiting the complexity of test cases submitted by the developers who are responsible for both test writing and data generation. Most of the test queries are simple single-table selection or aggregation queries with a handful of predicates.
II-B NL2SQL
NL2SQL is a more advanced service that translates natural language queries into complex SQL statements. It supports highly complex table schema with foreign key relationships; the generated SQL queries involving joins, aggregations, and nested subqueries over multiple tables. NL2SQL testing also requires the test data to be semantically coherent with the natural language question and the SQL. Due to privacy constraints and the scale of production data, NL2SQL testing relies on sampling from a restricted set of domains and tables. This limited production sampling significantly hampers test coverage and the ability to thoroughly evaluate across diverse scenarios. More specifically, the service has accumulated a large set of test SQL queries that are submitted by the internal developers, crowdsourcing and also provided by the customers with golden queries.
1Question: Which fake_devices have tasks related to fake_task_name with a start date after 2020-06-01, and what are the tasks and their start dates for the fake_devices? 2---------------------------------------------------- 3 4SELECT 5 name, 6 ARRAY_AGG(STRUCT(task, start_date)) 7FROM ( 8 SELECT 9 d.name, 10 t.name AS task, 11 FORMAT_TIMESTAMP("%Y-%m-%d", PARSE_TIMESTAMP("%Y-%m-%dT%H:%M:%E*S%Ez", t.start_date)) AS start_date 12 FROM fake_project_tasks AS t 13 INNER JOIN fake_devices AS d 14 ON t.project_id = d.id 15 WHERE ( 16 t.name LIKE ’%fake_task_name’ 17 ) 18 AND t.start_date > "2020-06-01" 19 ORDER BY start_date 20) 21GROUP BY name;
Figure 2: A natural language question and the golden SQL query with a join within the subquery.
Figure 2 illustrates a complex test case that require evaluating a golden SQL query with a join in the subquery. Some of the columns in the schema are also nested (proto and array types). Production sampling (graph mining-based) tends to focus on getting the most diverse set of instances, which may not guarantee a test data sample that is relevant for the given question. Given relatively small sample sizes, many of the test queries result in result sets with fewer rows than expected – yielding unrepresentative evaluation results. In our case, we generate test data that adheres to the arbitrarily complex schema structures and also relevant to the given test SQL structures using LLM. It is important to note that our generation process takes into account the semantics of the question and the schema. For instance, the LLM can make connection between the fake project tasks in the question and the fake device related task’s start date (t.start_date) to build up the context for relevant data generation. We observe that test data generated with the right semantic context can actually improve the test coverage significantly compared to production samples.
II-C Problem Definition
Let $\Lambda$ be a schema definition, like table schema or proto, with various types of fields/columns $\Xi$. The column set $\Xi$ includes primitive type (e.g., numerical, string, boolean) $\xi^{prim}$, categorical type $\xi^{cat}$ (e.g., enum), and nested type (e.g., struct, array, proto) $\xi^{nest}$ . $\xi^{nest}$ can be another (nested) schema definition $\Lambda_{i}$ that can have more nested structures within $\xi^{nest}_{i}$. Our work focuses on generating arbitrarily complex test data with any number of columns and with deeply nested structures, $\Xi=\{\xi^{prim}_{1},\xi^{cat}_{2},\xi^{nest}_{3},...,\xi_{n}\}$, for testing test SQL queries $\Omega=\{\omega_{1},\omega_{2},...,\omega_{m}\}$ and also without any training tabular dataset $\tau_{\Lambda}$ that confirms to $\Lambda$ – restricted production data access in real-world SQL code generation services.
Specifically, we leverage a pre-trained LLM, gemini, to generate high-fidelity and complex $\tau_{\Lambda}$ that enables testing $\Omega$. Table II illustrates the evaluation of test data generation for a large complex schema. Using the out-of-the-box gemini model (Vanilla LLM) for mock test data generation with $\Lambda$ resulted in incomplete data. Our generation method, incorporating strategic pre- and post-processing steps, results in the most complete test data, better than the manually prepared one by a human expert in an hour. Production sampling is not used, since the table is restricted. While the traditional manual data generation is very time-consuming, the LLM based generation runs much faster. We further discuss the per-query test data generation time and cost in Section VI-A.
III Test data Generation for SQL code generation services
Figure 3: The proposed LLM (gemini) based test data generation process. User provides context map (questions, schema, additional criteria) and test body (SQL query) as input. This data generation request triggers a series of data processing steps and modules. The goal is to generate syntactically correct and semantically relevant test data for SQL code generation services, like NL2SQL. We leverage the latest LLM capabilities to do this efficiently at scale.
To automate the generation of test data that aligns with user-specified criteria, we propose a novel test data generation pipeline leveraging the power of LLMs (gemini). Figure 3 provides a high-level overview of this process, illustrating the flow of data and processing from user input to the final validated mock test data output. The process begins with the user providing a context map – containing questions, schema definitions, and additional criteria—along with a test body, such as an SQL query or SQL function implementation. This input triggers a series of modules within the data generation library, including a pre-processor, context extractor, prompt builder, post-processor, and data validator. Each module plays a crucial role in refining and validating the generated data, ensuring it meets both the realized criteria from the schema and the query, as well as the user provided criteria.
III-A Pre-processor
This module will seed the signals needed to generate meaningful data, automatically based on information from the context map and the test SQL query body. Additionally this can also be passed from the user and any user provided input will override auto-generated ones. The signal will be a map of column name to column value (literal value, range or values, or a generator like incremental or other distribution, e.g., use a uniform distribution from 0 to 1 to populate this numeric column). These signals would be added to the context map and used later for post-processing.
Signals for Data Generation: Additionally the following signals would be needed to generate meaningful data:
  • Number of rows required.
  • Constraints in data points - e.g. specifying a currency in the example provided.
  • Column correlations.
  • Variations in data points - e.g. any field that needs to be different in each data point like customer ID.
  • SQL query and question (if applicable)
[…]
Figure 4: Pre-processor parses out subqueries (middle) from the original SQL query (top) for seeding and independent and dependent data generation. Join information is extracted to highlight the table and column association for the join (bottom).
Handling Subqueries: The pre-processor module treats each subquery as a separate query and associates each of the signals for data generation mentioned above separately for each subquery and the main query, as illustrated in Figure 4.
Handling Joins: Joins information is extracted from the SQL query separately to show table and column association for the join. The example SQL above the join information will be extracted as in Figure 4. Our method generates data (rows) per schema given a SQL query, the extracted join information will then be used in the post-processing to match the values in the table and columns present in the Join Constraints data structure. It is not possible to capture these nuances in a naive approach using Vanilla LLM.
III-B Context extractor
This module generates context relevant to the schema. We use a separate system to collect human written comments for a schema and proto from comments in data schema, source code and documentations. These comments are available at the schema/proto level and also individual column level. We use LLM to make these comments consistent and fill in missing columns. For protos, by themselves or as a column in a schema, we generate the proto description with comments for additional context for data generation. This component will also be able to extract column correlations from the SQL query body and annotated schema that would be useful in generating meaningful data.
Minimizing edge cases with constraints: We pass constraints extracted directly from the test SQL query, as illustrated in Figure 4, so that the model can explore and generate different values a relevant column in the predicates should take for the test (for instance, the LLM understands that it should generate for currency $\in\{``USD",``GBP",``EUR$"\} from the query in Figure 1). The same set of constraints are later used for the semantic quality verification of the generated data as well. In Section VI-B we discuss another example where our LLM-based generation successfully generate test data around date_trunc(logdate,quarter)date\_trunc(logdate,quarter) constraint, leaving no quarter without test data.
III-C Prompt builder
The test data generation prompt is constructed using the following template in Figure 5. The template contains some system instructions to specify the task and also to prevent common hallucination problems, like repeating the same column more than once, carefully validating and matching the column/field names to the schema.
You are a GoogleSQL expert. GoogleSQL is a Google internal implementation of the SQL query language. Given a proto description, generate sample data in textproto format.
Use only the proto names present in the proto files. Do not create columns that do not exist. Check if the columns match with the proto description.
Instructions for generating sample data:
Pay careful attention to the datatypes in the proto columns. Don’t make up the proto field names. Carefully match the proto field names to their data types. Don’t repeat the fields more than once. Ignore any deprecated fields in the proto. Generate the nested proto fields in the proto file. Instructions for specific columns in the given schema: {constraints}
Here are the steps that each question should follow:
1. Generate the proto column name only if present in the proto file. 2. If you have already generated data for a field, please move to the next field. 3. Generate non-zero values for every numeric field. ”””
system_instruction = f”””
{system_prompt_prefix}
{data_generation_signals}
user_instruction = f”””
Task: Given the following proto description, generate
{number_of_data_points} rows of sample data for col_names columns.
{user_input}
Proto/schema Description:
{proto_description}
mock_data_gen_prompt =
f”””
system:{system_instruction}
user:{user_instruction}
model:
Figure 5: Data generation prompt template with system and general instructions. Custom generation seeding instructions can be passed as part of data_generation_signals and user_instruction.
1---------------------------------------------------- 2# Without column grouping: 3 4json[{ 5 "contract_created_date": "2023-08-15", 6 "contract_created_month": "2023-08" 7 "contract_created_year": 2023, 8 "contract_effective_date": "2023-04-01", 9 "contract_closed_date": "2023-03-30", 10 "contract_executed_date": "2023-03-29", 11 "contract_expiry_date": "2024-03-31", 12 "fully_signed_time": "2023-03-29T10:00:00Z", "contract_termination_date": "2023-12-31", 13}, 14...] 15---------------------------------------------------- 16# With column grouping: 17 18json[{ 19 "contract_created_date": "2023-10-27", 20 "contract_created_month": 10, 21 "contract_created_year": 2023, 22 "contract_effective_date": "2023-11-01", 23 "contract_closed_date": "2024-10-31", 24 "contract_executed_date": "2023-10-30", 25 "contract_expiry_date": "2025-10-31", 26 "fully_signed_time": "2023-10-30T10:00:00Z", 27 "contract_termination_date": "2024-05-15", 28 "contract_termination_date_option": "SpecificDate", 29 "contract_months_from_effective_date": 12, 30 "simba_contract_create_date": "2023-10-27", 31}, 32...]
Figure 6: Column grouping based on column correlation helps generating more coherent set of values. Without column grouping (above) the contract effective (start) date is before the contract created date; with column grouping (below) all dates are aligned and correct.
1---------------------------------------------------- 2# column annotation 3 4"terminated_by_clm_system" // "Ifthecontractwasterminated,thisfieldindicatestheContractLifecycleManagementSystemofthecontract(s)responsibleforitstermination.Thisfieldispartofthesourceinformationaboutthiscontract’slineagewithintheContractFamilyLinkage." 5 6---------------------------------------------------- 7# Generated column value without annotation: 8"terminated_by_clm_system": "TerminationSystem", 9 10---------------------------------------------------- 11# Generated column value with annotation: 12"terminated_by_clm_system": "InternalCLMSystemCode",
Figure 7: An example column annotation prepared by the context extractor from human written comments, source code and documentation. This allows the generation to look up and consider valid set of column values, instead of random (hallucinated) strings.
When we make the LLM call for generation, we make the request for each nested field separately on a separate thread (see Section VI-A for more details). For the primitive/scalar fields, we group and pass one generation call to handle 30 fields at a time. The grouping is based on the column correlations identified by the context extractor. This helps maintain semantic consistency across the correlated columns, as supposed to generating a valid value for each column individually (see Figure 6). The column grouping sometimes leads to unexpected hallucinations where the LLM introduces fake (non-existent) columns, especially when the schema is small with a fewer number of primitive fields (see Section VI-C for an example). To guard against such hallucination, we also keep track of the requested column generations so that any hallucinated columns can be detected and removed. The schema (proto description) passed here also contains column/field annotations prepared by the context extractor. The annotations are attached to the corresponding schema elements as comments. This also improves the semantics of the generated data, as illustrated in Figure 7.
III-D Post-processor
This module can add the data constraints and variations by copying the basic data generated by the LLM or existing data written by the user. The extracted join constraints in the pre-pro in the pre-processing are also enforced by the post-processor, where the column values in the joined tables at the join columns are matched and copied from one table data to another. Since the constraints and variations will affect only a smaller number of columns in a proto, we could easily just modify those components.
This module will not make any LLM calls and will update the relevant fields identified in the preprocessor module. This will also provide us a chance to correct any problems with the generated data. This will also reduce the total time taken for data generation as the LLM call latency scales linearly with the number of data points to be generated. This will also make the changes more deterministic in nature.
III-E Data validator
We would evaluate the generated data using a proto parser [12]. This has an in-built check for the datatype and any other constraint on the proto fields. In case of any failure, we would be retrying the data generation, at most 2 retires, along with an error message. The integrated testing will be done by the framework.
TABLE III: Enforced data validation rules. The generator extracts most of the syntactic and semantic constraints automatically, including the join paths. Additional input signals used for generation are also used for validation, like column correlation and constraints.
Rule Validation Requirements Validation Approach
$r_{1}$ No missing columns Use a proto parser to check for missing columns or proto fields.
$r_{2}$ Valid data types Use a proto parser to check for invalid data types. A schema column or proto field is invalid if its data type is different from the one specified in the fetched schema/proto.
$r_{3}$ Grouped column correlation Use statistical tools, like Pearson Correlation (numerical vs. numerical) and Chi-Square Test (categorical vs. categorical), to assert non-zero column correlations.
$r_{4}$ Join constraints Matches any join constraints extracted in the pre-processing. The column values in the joined tables at the join columns would be matched/copied from one table data to another.
$r_{5}$ Semantic constraints Semantic constraints are extracted from the test SQL query and they are validated using LLM-as-a-judge.
You are a GoogleSQL expert. GoogleSQL is a Google internal implementation of the SQL query language. Given data in json format and constrains for particular columns in data, validate the data.
Instructions for validation: Check the values required for a column under the constraints section. Check for the same column in the data section. If the column is not present, please return NOT PRESENT. If the columns is present, please check if the data is present in the constraint for that column. If point 4 is satisfied, please print VALID. If point 4 is not satisfied, please print NOT VALID. user_prompt = f”””
Task: Given the following Data and Constraints, validate the data and return one of VALID, NOT VALID, NOT PRESENT.
Data:
{data}
Constraints:
{constraints}
Figure 8: Semantic validation prompt template for generated test data. We also pass in the extracted SQL constraints, including the join information and semantic constraints (e.g., relevant column values given the test query and question)
Semantic constraints and validation is not easy to capture with conventional approaches, and we use the LLM for semantic validation (LLM-as-a-judge [33]) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 8. Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks.
IV Evaluation results
In this section, we present evaluation results of our proposed LLM-based test data generator. The evaluation investigates three key aspects that highlight the efficacy and advantages of our approach. First, we assess the test coverage achieved by the generated data, focusing on its ability to facilitate more complex SQL query testing. Second, we evaluate the completeness and correctness of the generation for complex nested data structures – as commonly encountered in our SQL code generation workloads. This evaluation is based on the criteria used in Table 2, ensuring a quantitative assessment of the generated data’s structural integrity and fidelity to the defined schemas and protos. Through these evaluations, we demonstrate the superiority of our approach in creating high-quality, comprehensive test datasets. Lastly, we evaluate the effectiveness of our data generation methodology in comparison to production data sampling. Based on a limited set of authorized production tables from restricted domains, we perform a comparative analysis of the generated data’s ability to provide meaningful test data with respect to the target test SQL query/function, covering all the edge cases for semantically relevant predicate columns.
IV-A Evaluation setup
Test cases are submitted by the developers who work on the code generation services. In SQL Code Assistant, a developer may generate SQL functions using the assistant service for testing GoogleSQL modules and use the generated SQL functions with some test fixtures (e.g., manually curated schema/proto instances – often incomplete and neglect nested fields). We sampled 20 SQL Code Assistant generated test cases that require querying over complex data structures – out of 129 disabled (buildable, but unexecutable) test cases. In NL2SQL, each submitted test case contains a natural language question and a golden GoogleSQL query, verified by engineers expert in GoogleSQL. We sampled 148 test cases out of which, only 45 had authorized access to production tables for testing.
For our test data generation pipeline, we used the latest gemini-flash-2.0 model and with 10 concurrent threads handling nested structures (e.g., nested field that leads to another schema/proto structure). The threading happens at the same level of nesting to ensure that there is no data dependency between concurrent threads.
IV-B Test coverage with more complex SQL query testing
We assess the test coverage achieved by the generated test data. Our generation method can generate test data for complex schema with deeply nested structures. This facilitates more comprehensive testing, enabling more complex SQL test cases that were disabled due to the lack of such complex test data. Table IV presents improvement (%) in test coverage and test accuracy – we report the improvement instead of the raw test coverage values. The improvement in NL2SQL testing coverage is more significant at $69.59\%$ because many available test cases were disabled due to restricted production data access. Notice that the generated test data also improves NL2SQL test accuracy. To our surprise, there were many cases where manually curated production test data that failed to differentiate predicted and golden SQL queries. For instance, some aggregate queries with wrong predicted filter predicates resulted in the correct answers (false positives); many test queries with INNER JOIN were flagged as incorrect (false negatives) against the manually curated test data, whereas adding sufficient test data for all the edge cases for the relevant columns used in the SQL query (Section VI-B) resulted in correct matching results. In SQL Code Assistant there are a lot more disabled test cases (mostly disabled because it is challenging to generate complex nested test data instances manually), and we expect to roll out new generated test data for those in the near future, which should increase the coverage further.
TABLE IV: Test coverage and test accuracy changes (%) enabled by complex data generation and with more complex SQL test cases. NL2SQL test coverage improvement is much significant because many complex test cases were disabled due to restricted production table access before; NL2SQL test accuracy also improved significantly, since the manually curated test data allowed many false positives and false negatives.
Metric SQL Code Assistant NL2SQL
Test coverage (%) +12.58 +69.59
Test accuracy (%) +0.0 +20.4
IV-C Generation quality evaluation with complex nested data structures
Next, we evaluate the completeness and correctness of the generation for complex nested data structures. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types.
Figure 9: Structural integrity & adherence to complex schema analysis. The scores are normalized (0 to 1) and the higher the better. We noticed some nested schema/proto definitions are inaccessible due to staleness and removal (invalid proto). Our data generation can generate test data compliant to complex nested schema structures from real-world scenarios.
Figure 9 presents our structural integrity & adherence to complex data schema evaluation results. We use a subset of key evaluation criteria derived from the ones used in Table II – we excluded the criteria for primary key generation, which is always true for our test data. The evaluation scores are normalized (0 to 1) and the higher the better. Overall, our generated test data adheres very well to complex schema with deeply nested column types encountered in production SQL code generation services. There are some cases where nested field (proto) definitions are stale or removed, in which case the LLM guesses the field names and types for the inaccessible nested proto structures. Note that the scores are much higher (and near perfect) without such invalid proto definition cases w/o invalid proto. Note that the correct nesting levels scores for both with and without invalid proto are the same because we skip over those invalid nested structures in the aggregation – since we cannot assess the correct nesting levels for unknown proto. We typically see 3-level of nesting in complex SQL testing scenarios, and as deep as 5 levels.
IV-D Semantic data constraints validation with LLM-as-a-judge
The generator extracts semantic data constraints, like which column values are relevant to the given SQL query in the pre-processing step. We use the LLM as LLM-as-a-judge to validate the generated test data against the extracted semantic constraints. If some generated data instances fail to satisfy the constraints according to the LLM, then we retry with an additional hint that the previous generation failed to satisfy the failed semantic constraints. In all test cases, we have successfully generated test data instances after 1 or 2 retries. Interestingly, we observed that the LLM can hallucinate and actually pass invalid data instances as valid. Table V illustrates how generating multiple data instances per test case can limit the hallucination error rates, defined as the proportion of test cases with no semantically valid test data instances after $n$ instances generated. We report the hallucination error rates for NL2SQL test cases only because the unit test cases for SQL Code Assistant are too simple and contains simple semantic constraints, like COUNTIF(experiment_group = “GROUP A”) AS group_a_count. While most semantic constraint violations are resolved with a single additional data generation attempt ($n=2$), a few constraints remain challenging for the LLM-as-a-judge to validate. Specifically, complex datetime filter constraints that require type casting, time zone conversion, and date extraction consistently failed even after five attempts ($n=5$). For example, consider the constraint DATE(TIMESTAMP_SECONDS(CAST(status.time_processed _sec AS INT64)), ’America/Los_Angeles’) $\geq$ “2023-07-01”. Given a generated value like ”1678886400” for time_processed_sec, the validator correctly performed the initial conversion, recognizing that the timestamp corresponds to the date “2023-03-15”. However, it then failed to perform the final step of evaluating whether this date satisfied the constraint (i.e., that “2023-03-15” is not greater than or equal to “2023-07-01”). A potential remedy involves invoking a reasoning model or chain-of-thought prompting [34], though this significantly increases generation time. We plan to explore more cost-efficient ways to address these hallucination errors in future work.
TABLE V: LLM hallucination error rates for semantic data validation with repeated data generation ($n$, the number of test data instances) per test case. By generating multiple instances $n>1$, we get at least one test instance that satisfies the semantic constraints for most of the test cases. Semantic validation for SQL Code Assistant test cases was perfect, since the unit test cases did not involve any complex semantic constraints.
Number of instances per test ($n$)
1 2 3 4 5
NL2SQL 6.08% 2.02% 2.02% 2.02% 2.02%
IV-E A comparative study VS. production sample
Here we evaluate the effectiveness of our data generation methodology in comparison to production data sampling. We use one production table authorized for testing to generate data samples of different sizes. To compare the quality of our generated test data against the production samples, we compare the number of distinct values of the relevant predicates/columns extracted from the question and golden SQL query. We use a test case submitted by an anonymous developer, “How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom(GB) during Q1 2024?,” for the test production table. Note that evaluating the NL2SQL output result is not meaningful as it should return different answers depending on different samples. There are two relevant columns also used as predicates, describing different country codes and semantic tags for the Shorts videos.
(a) Country codes
(b) Semantic tags
Figure 10: The number of distinct values generated for relevant predicates columns in the test SQL query and the natural language question, ”How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom (GB) during Q1 2024?”
Figure 10 presents the number of distinct values for the (a) country codes column and (b) semantic tags column. Because the question is asking for the (identified as) female case, we also track the counts by (identified as) male and (identified as) female. In case of country codes, the production samples Prod include a more diverse set of countries than the Mock ones. This is OK because the question is looking for the answer in the United Kingdom (GB) specifically – and it is likely that our generation is less diverse because of this. At the same time, we also notice an imbalanced sampling between Prod/M and Prod/F in the production samples, which is not desirable as the question seeks to answer for the (identified as) female case. In case of semantic tags, which can entail different genres and keywords for different gaming contents (FPS, MMO, RPG, tutorial, review, etc.), our generated test data Mock does include a more diverse set of values. In both cases, the data is well balanced between the two select genders. Note that the test data is generated with syntax and semantic validations as described in Section III-E, but the values can still be unrealistic as it is not learned from the actual production data (not available). Our generation approach also takes the test query/function into account to generate data that is more semantically aligned to the query. We discuss another example in Section VI-B.
V Related works
Test data generation for complex SQL testing: Test data generation is crucial for SQL query validation, query rewriting, automated query grading, and NL2SQL tasks. Common strategies involve using manually curated test data, production data sampling, constraints based data generation for table schema and SQL query [5, 1, 3, 4, 6]. Production data sampling can provide the most realistic test data for any complex schema, but it is often prohibited or infeasible in industrial contexts where service providers lack access to user data due to stringent privacy and security protections/policies. One of the key challenges in generating realistic high-fidelity test data is handling more complex nested schema structures and their inter-dependencies. Many SQL equivalence testing and constraints based SQL test data generation do not support arbitrarily nested structures and nested fields (array, struct and/or proto types) [11, 5] – limiting the test use cases to less complex SQL queries.
Model-based tabular data generation: Despite the emerging interest in LLMs for test data generation [16, 19, 20, 21], existing approaches often fail to fully address the complex testing requirements of industrial applications. While LLMs can provide realistic textual data and synthesize tabular data that adheres to schema format [27], the out-of-the-box LLMs fall short at capturing subtle nuances and relationships betwen data atrributes [25]. We also found that the vanilla LLM generation becomes more challenging and incomplete for a large nested schema (Table II). Other deep generative model-based techniques have been used for synthetic tabular data [29, 30, 31]. While producing realistic data examples that closely approximate the underlying distributions, the traditional tabular data generation solutions focus on flat schema without nested fields (struct, array, proto). Another challenge is to source good training datasets for learning the tabular generation model, since the production data is often restricted in the real-world testing environments. In this work, we propose a complex test data generation method based on a pre-trained LLM, gemini, and without any post-training.
VI Discussion
VI-A Scaling and cost of the generation
While much more efficient and scalable than manual test data generation, as illustrated in Table II, our test generation process can easily be deployed and parallelized across different batch of test queries/schemas, and also across different nested structures/fields within the same schema. The test data generation efficiency is important, also the main pain-point shared by the practitioners (e.g., developers submitting test queries and manual test data), as the data generation time takes much longer than the actual test generation/execution time.
TABLE VI: Average ($\pm$ 1-std) number of accumulated input & output tokens, generation time per test data generation for the SQL code generation services. One can estimate the actual monetary cost using a (external) pricing model for gemini-2.0-flash.
Metric SQL Code Assistant NL2SQL
Input tokens 162199.5 ($\pm$ 350852.1) 57082.3 ($\pm$ 102381.8)
Output tokens 3890 ($\pm$ 3121.5) 8054.2 ($\pm$ 13040.1)
Generation time (sec) 66.0 ($\pm$ 53.6) 57.5 ($\pm$ 64.36)
Table VI presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model [35] (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
VI-B Generating for test SQL query semantics
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query [36, 37]. While the NL2SQL prediction quality evaluation is beyond the scope of this study, a meaningful test data generation should provide a set of test data instances that are semantically relevant to the target SQL query. One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section IV-E.
1SELECT date_trunc(logdate, quarter), sum(conversions) / sum(clicks) 2FROM fake_table_1 3WHERE result_type = ’TEXT_AD’ AND logdate BETWEEN ’2022-01-01’ AND ’2022-12-31’ 4GROUP BY date_trunc(logdate, quarter);
Figure 11: Meaningful test data generation for this business analytics query requires generating data instances for each quarter in year 2022.
To illustrate the problem, consider a business analytic query in Figure 11. A meaningful test data for this query should ensure that enough example conversions instances are present across all the quarters in the requested date range. Our LLM-based generator can recognize such semantic importance of date_trunc(logdate, quater) and generate data for all quarters in the range. Moreover, if one has a domain knowledge about the data, such as “conversions were better in Q4 than Q1,” then it can be passed to the generator as user_criteria as part of input, which can guide the generation to favor Q4 instances. Such understanding and reasoning capabilities of LLMs can scale semantic-aware test data generation.
VI-C Generation error analysis
We observe some common error patterns in our data generation process:
  • Output data formatting error. We request output in JSON format, but the actual responses had incorrect JSON formatting failing the JSON parser. This can be fixed with retries.
  • Missing fields from large and complex schema, which requires some post-processing steps to identify and retry.
  • Generating fewer than requested amount of data/rows, which can be fixed with retries to generate data incrementally.
  • For some schemas with a very small number of primitive columns, the LLM hallucinated and introduced new columns. Care is needed to keep track of requested columns/fields for each (nested) schema.
  • Missing some edge cases relevant to given SQL queries/functions.
1json[{ 2 "fake_service_id": 1234567890, 3 "user_dims": { 4 "dimension1": "value1", 5 "dimension2": 10 6 }, 7 "id": 12345, 8 "customer_dims": { 9 "dimension3": "value3", 10 "dimension4": 20 11 }, 12 "customer_details": { 13 "detail1": "detailValue1", 14 "detail2": 30 15 }, 16 "fake_service_dump_freshness_usec": 1678886400000000, 17 "user_details": { 18 "detail3": "detailValue3", 19 "detail4": 40 20 } 21 }, 22 ...]
Figure 12: LLM hallucinates to introduce new columns. Only “fake_service_id”, “fake_service_dump_freshness_usec”, and (user) “id” are requested, but new fields are being introduced in the data.
We have implemented some number of safeguards against LLM hallucination. Most notably, we have observed in some simple schema cases, the LLM introduced new fields – even nested ones. Figure 12 illustrates one such example where many more new fields are introduced next to just 3 required fields; the LLM even introduces nested fields.
The last item is very difficult to fully address. We use LLM-as-a-judge to validate the semantic coherence of generated data, which runs the risk of false positives, but we do not have an oracle (or access to the real production tables) to ensure all relevant values are covered and generated. If we can provide some domain knowledge, as discussed in Section VI-B, that could help inform the LLM to generate for particular cases. This is an important problem to further improve the fidelity of mock data generation in the production settings. There are also techniques to statistically impute missing values and estimate potential impact of missing values on query results [38, 39], which would be interesting to consider in the test data generation settings.
VII conclusion and future work
In this paper, we addressed the critical challenge of generating high-fidelity and complex test data for Google SQL code generation services, particularly in environments where production data access is restricted. We highlighted the limitations of traditional data generation methods, focusing on their inability to handle large, deeply nested schemas and maintain semantic integrity to the target test cases. By leveraging the latest gemini-2.0-flash model, coupled with strategic pre- and post-processing, we demonstrated the capability to generate comprehensive test data instances compliant to large complex schemas that effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries. Our experiments demonstrate the practical effectiveness of our approach, delivering a reliable method for generating realistic test data essential in industrial environments where access to production datasets is restricted. While our implementation targeted Google SQL, the fundamental methodology is dialect-agnostic and should be readily applicable to other SQL code generation services. By deploying our test data generator, we’ve enabled more thorough SQL code testing, resulting in significantly improved test coverage.
As part of our future work, we plan to extend the capabilities of our framework to support other nested data types, such as XML, JSON, and Open API schemas. This will enable us to address a broader range of real-world applications beyond SQL code generation services, including RESTful API testing. Furthermore, we aim to explore techniques for enhancing the semantic coherence of generated data, ensuring that it not only adheres to structural constraints but also accurately reflects real-world business context and semantics. By expanding the scope and capabilities of our approach, we seek to provide a comprehensive solution for generating high-quality test data that addresses the evolving needs of industrial and large-scale data service development and testing.
Acknowledgment
We would like to thank John Cieslewicz, Kashmira Phalak, Yannis Papakonstantinou and the anonymous reviewers for their valuable feedback. We also thank Yuchen Zhang for his contribution in the early stages of this project.